OpenStudy (anonymous):

Proof for fundamental theorem for line integral? I don't get how to use chain rule to proof this theorem

OpenStudy (anonymous):

\[\int\limits_{b}^{a} \frac{ \partial f }{ \partial x } \frac{ dx }{dt } +\frac{ \partial f }{ \partial y } \frac{ dy }{dt }+\frac{ \partial f }{ \partial z } \frac{ dz }{dt } dt\]

OpenStudy (anonymous):

that's what I got up to, but I don't get how they used chain rule to get

OpenStudy (anonymous):

integration of d/dt (f(rt)) dt from b to a

OpenStudy (anonymous):

that integral above was suppose to be from b to a

OpenStudy (anonymous):

That's the site I'm using I don't how he just used chain rule though

OpenStudy (phi):

Did you look at his explanation for the chain rule http://tutorial.math.lamar.edu/Classes/CalcIII/ChainRule.aspx

OpenStudy (anonymous):

yes, that example uses x so it makes sense that dx/dx cancels but in this case its in terms of dt

OpenStudy (anonymous):

Oh nvm I think I reasoned it out

OpenStudy (anonymous):

thanks for our help